OpenStudy (insa):

i just need the equation for this question: one third of the sum of two angles is 60 degree and one quarter of their difference is 28 degree. find the two angles.

OpenStudy (ajprincess):

Let the angles be x and y. sum=x+y (x+y)/3=60 difference=x-y (x-y)/4=28

OpenStudy (insa):

ths alot @ajprincess

OpenStudy (ajprincess):

u r welcome:)